a car travels 52 miles per hour for 6 minutes, 46 miles per hour for 3 minutes, 40 miles per hour for 15 minutes and 44 miles per hour for the remaining time during a 30-minute journey. What is the average speed of his car?
1
Expert's answer
2020-10-09T07:11:42-0400

The average speed is total distance traveled divided by elapsed time. We will convert minutes to hours by division by 60 (try to think why).

Thus, all we must do is to calculate each piece of distance traveled with certain speed during certain time, add these pieces altogether, and divide this by elapsed time.


"s=\frac{D}{T}=\\\space\\
=\frac{52\cdot6/60+46\cdot3/60+40\cdot15/60+44(30-6-3-15)/60}{30/60}=\\\space\\
=43.8\text{ mph}"